OpenStudy (anonymous):

Partial Differentiation Problem: Determine fx when f(x, y) = (2x + y)/ (x + 2y) @agent0smith

OpenStudy (anonymous):

Is: \[f(x)=\frac{2x+y}{x+2y}\]

OpenStudy (anonymous):

yes @KeithAfasCalcLover

OpenStudy (zarkon):

just use the quotient rule

OpenStudy (anonymous):

So you would have: \[\frac{\partial}{\partial x}\frac{2x+y}{x+2y}=\frac{\partial}{\partial x}\frac{f(x,y)}{g(x,y)}\] So we can obtain: \[\frac{\partial}{\partial x}\frac{f(x,y)}{g(x,y)}=\frac{f_x(x,y)g(x,y)-g_x(x,y)f(x,y)}{g(x,y)^2}\]

OpenStudy (anonymous):

Clear so far?

OpenStudy (zarkon):

you probably shouldn't use f(x,y) in your explanation since it is already defined to be the entire original problem.

OpenStudy (anonymous):

Fair point. But I was just stating in general that for any \(\frac{f(x,y)}{g(x,y)}\), that would be the derivative

OpenStudy (zarkon):

I know...just wanted to avoid confusion for the OP

OpenStudy (anonymous):

What is OP?

OpenStudy (zarkon):

original post (or poster)

OpenStudy (anonymous):

Will avoid that next time. Good Point!

OpenStudy (zarkon):

hopefully pancakeslover can add something here since he/she now has a formula

OpenStudy (anonymous):

I agree! I like to give a general solution first. But I must go haha so I will talk to you later! Good Night everybody!

OpenStudy (anonymous):

no i got it!! thank you! I just have to keep y as a constant and differentiate with respect to x. Thank you!